Derive 4 th kinematics equation of motion. …...............

Derive 4th kinematics equation of motion.                                                                                 …...............

Grade:11

1 Answers

mohan
6 years ago
It is sn-sn-1 =(un+1/2an^2)-(u(n-1)+1/2a(n-1) ^2)So fourth equation of kinematics is=u+a(n-1/2) For u=0 the equation becomes a(n-1/2a)